Its Simultaneous equations.
Please solve ;
x-y=4
4x-6y=21
Thanks !

Answer by shweta(56)

Given:
x-y =4 ......(1)
4x-6y=21 ......(2)
these are the equations with two variable 'x' and 'y'. First we need to remove any one variable.This can be done when the coefficient of any variable is same so that it get cancelled. Lets take variable'x'
We multiply first equation with 4 because the coefficient of 'x' in the second equation is 4
( x-y= 4 ) *4
4x-4y = 16 ......(3)
Now consider equation 2 and 3
4x-6y =21
4x-4y =16
In order to cancel the alike term ,we will subtract both the equations
4x-6y =21
4x-4y =16
- + -
0-2y = 5 { 4x-4y= 0,(-6y)+4y= -2y , 21-16 =5 }
-2y = 5
y = -5/2 .....(4)
Now substitute the value of y in ,say equation 1
x-y =4
x-(-5/2) = 4
x+ 5/2 =4
Subtracting 5/2 from both the sides to get the value of x
x = 4-5/2
x =3/2
Hence, solved